# Quantum Mechanics Relative to a Quantum Reference System: a Relative State Approach

### v2

@@ VII Comparison with Some Past Works
Of course, as a quantum reference frame theory, the philosophical interpretation of quantum mechanics presented in this paper is much closer to that of the relational quantum mechanics (Rovelli:1995fv, ; Rovelli:2021thu, ), and has little to do with the “relative state” in the so-called many-worlds interpretation, despite the semantic similarity between “relational” and “relative”—we do not wish to get bogged down in semantic quibbles here, and we also don’t want to get entangled in the philosophical disputes among various schools of thought derived from the relational interpretation. The fundamental viewpoint we advocate in this paper is that quantum states are relative (w.r.t. quantum apparatus), and physical properties can only be defined relatively. Hence, there is no global section (of the non-trivial fiber bundle) serving as a global wave function and moreover no global underlying “hidden variable”. The “relative state” is a local section on a non-trivial fiber bundle, whereas the many-worlds interpretation still posits a global wave function, with observers merely perceiving branches of this global wave function. Of course, unlike relational quantum mechanics, which merely offers a new interpretation of the existing mathematical formalism of quantum mechanics, we strive to more thoroughly incorporate the spirit of general relativity into the framework of quantum mechanics.
@@ VIII Conclusions
(3) The entangled state of the under-study state $|X\rangle$ and the clock apparatus state $|T\rangle$, prepared during calibration, does not evolve with respect to external parameters. However, the relative evolution of the under-study quantum state $|X\rangle$ relative to the clock state $|T\rangle$ in the intrinsic quantum theory can be described by a evolution equation of the total quantum state space coming from the Ricci-flat Kähler-Einstein equation. The evolution equation is a generally covariant equation that no longer refers to any external absolute parameters and constrained by a vanishing Hamiltonian, but only relies on internal relative descriptions. In this sense, this framework achieves a geometric formulation of a quantum theory describing the under-study quantum system and the quantum apparatus. This equation has a linear and non-relativistic approximation to the Schrödinger equation.
@@
(7) The relative evolution equation for the quantum state of the one-dimensional object’s position relative to a single-degree-of-freedom quantum clock reference system has already been a high dimensional evolution equation. Therefore, it can be imagined that if more degrees of freedom of the quantum spacetime reference system are involved, this canonical framework will become even more complicated. This canonical framework, based on differential equations for quantum state evolution, is more suitable for conceptually demonstrating its relationship with the standard quantum mechanical Schrödinger equation and its probability interpretation, helping us observe more effects that are not easily seen through the functional integral approach. On the other hand, the action and functional integral quantization framework is actually more amenable to practical mathematical calculations and to study quantum gravity. Each method has its own advantages and disadvantages. To facilitate comparison between these two methods, we will provide the functional method for a single-degree-of-freedom quantum clock in the Appendix for easy comparison. For the functional treatment of quantum spacetime reference systems with more degrees of freedom, please refer to (Luo:2021zpi, ; 2024arXiv240809630L, ; 2024chinarxiv, ).
@@ Acknowledgements.
When this model of a quantum clock with a single degree of freedom is generalized to a quantum spacetime reference frame with multiple degrees of freedom, the aforementioned results are extended in a similar manner: the action $S_{T}$ of the single-degree-of-freedom quantum clock is generalized to a non-linear sigma model with 4-degrees of freedom $X,Y,Z,T$, serving as the quantum spacetime reference frame fields. Meanwhile, the action $S_{X}$ of the object’s position can be generalized to the action of a general quantum field, e.g. $\phi$, representing the quantum system under-study. The global parameter $\tau$ is extended from a single parameter to the four Minkowski coordinate parameters $x,y,z,t$ of the laboratory spacetime, which act as the base space of the non-linear sigma model, while the target space corresponds to the actual spacetime $X,Y,Z,T$. The resulting model, under the semiclassical approximation, yields the effective action of the quantum field $\phi$ on a general curved spacetime background. Furthermore, when considering the effects of second-order quantum fluctuations of the quantum spacetime reference frame fields beyond the semiclassical approximation, the effects of “inertial forces”—namely, gravity—automatically emerge, and, under certain approximation, it takes on the form of Einstein gravity (Luo:2021zpi, ; 2024arXiv240809630L, ; Luo:2025cer, ).

### v3

@@ VII Comparison with Some Past Works
Of course, as a quantum reference frame theory, the philosophical interpretation of quantum mechanics presented in this paper is much closer to that of the relational quantum mechanics (Rovelli:1995fv, ; Rovelli:2021thu, ), and has little to do with the “relative state” in the so-called many-worlds interpretation, despite the semantic similarity between “relational” and “relative”—we do not wish to get bogged down in semantic quibbles here, and we also do not want to get entangled in the philosophical disputes among various schools of thought derived from the relational interpretation. The fundamental viewpoint we advocate in this paper is that quantum states are relative (w.r.t. quantum apparatus), and physical properties can only be defined relatively. Hence, there is no global section (of the non-trivial fiber bundle) serving as a global wave function and moreover no global underlying “hidden variable”. The “relative state” is a local section on a non-trivial fiber bundle, whereas the many-worlds interpretation still posits a global wave function, with observers merely perceiving branches of this global wave function. Of course, unlike relational quantum mechanics, which merely offers a new interpretation of the existing mathematical formalism of quantum mechanics, we strive to more thoroughly incorporate the spirit of general relativity into the framework of quantum mechanics.
@@ VIII Conclusions
(3) The entangled state of the under-study state $|X\rangle$ and the clock apparatus state $|T\rangle$, prepared during calibration, does not evolve with respect to external parameters. However, the relative evolution of the under-study quantum state $|X\rangle$ relative to the clock state $|T\rangle$ in the intrinsic quantum theory can be described by an evolution equation of the total quantum state space coming from the Ricci-flat Kähler-Einstein equation. The evolution equation is a generally covariant equation that no longer refers to any external absolute parameters and constrained by a vanishing Hamiltonian, but only relies on internal relative descriptions. In this sense, this framework achieves a geometric formulation of a quantum theory describing the under-study quantum system relative to the quantum apparatus. This equation has a linear and non-relativistic approximation to the Schrödinger equation.
@@
(7) The relative evolution equation for the quantum state of the one-dimensional object’s position relative to a single-degree-of-freedom quantum clock reference system has already been a high dimensional evolution equation. Therefore, it can be imagined that if more degrees of freedom of the quantum spacetime reference system are involved, this canonical framework will become even more complicated. This canonical framework, based on differential equations for quantum state evolution, is more suitable for conceptually demonstrating its relationship with the standard quantum mechanical Schrödinger equation and its probability interpretation, helping us observe more effects that are not easily seen through the functional integral approach. On the other hand, the action and functional integral quantization framework is actually more amenable to practical mathematical calculations and to study quantum gravity. Each method has its own advantages and disadvantages. To facilitate comparison between these two methods, we will provide the functional method for a single-degree-of-freedom quantum clock in the Appendix for easy comparison. For the functional treatment of quantum spacetime reference systems with more degrees of freedom, please refer to (Luo:2021zpi, ; 2024arXiv240809630L, ; 2024chinaxiv, ).
@@ Acknowledgements.
When this model of a quantum clock with a single degree of freedom is generalized to a quantum spacetime reference frame with multiple degrees of freedom, the aforementioned results are extended in a similar manner: the action $S_{T}$ of the single-degree-of-freedom quantum clock is generalized to a non-linear sigma model with 4-degrees of freedom $X,Y,Z,T$, serving as the quantum spacetime reference frame fields having a vanished Hamiltonian. Meanwhile, the action $S_{X}$ of the object’s position can be generalized to the action of a general quantum field, e.g. $\phi$, representing the quantum system under-study. The global parameter $\tau$ is extended from a single parameter to the four Minkowski coordinate parameters $x,y,z,t$ of the laboratory spacetime, which act as the base space of the non-linear sigma model, while the target space corresponds to the actual spacetime $X,Y,Z,T$. The resulting model, under the semiclassical approximation, yields the effective action of the quantum field $\phi$ on a general curved spacetime background. Furthermore, when considering the effects of second-order quantum fluctuations of the quantum spacetime reference frame fields beyond the semiclassical approximation, the effects of “inertial forces”—namely, gravity—automatically emerge, and, under certain approximation, it takes on the form of Einstein gravity (Luo:2021zpi, ; 2024arXiv240809630L, ; Luo:2025cer, ).